OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $395. How many sophomores bought tickets? 55 75 90 95

OpenStudy (aroub):

95

OpenStudy (anonymous):

Use a system of equations to solve for s, letting s represent the number of sophomores and f represent freshmen. s+f=150 2f+3s=395

OpenStudy (aroub):

=> sophomores = x freshmen= 150-x => 2(150+x)+3x=395 300+x=395 x=95